# Uncertainty Quantification of Surrogate Models using Conformal Prediction

### v2

@@ C.2 Model and Training
We train U-Nets and FNOs to map the spatio-temporal evolution of the field variables. For the U-Nets, the network takes in the first 20 time instances (T_in) to map the next 30 time instances (step). The U-net performs a feed-forward mapping without any autoregressive roll-outs. For the FNO we deploy an auto-regressive structure that performs time rollouts allowing us to map the initial time steps in a recursive manner up until the desired time instance (T_out). Each model autoregressively models the evolution of the field variable up until the $80^{th}$ time instance. The U-Net architecture can be found in Table 6 and the FNO in Table 7. For the case of STD, the architecture is modified with 2D dropout layers following each encoder and decoder of the U-Net and after each Fourier layer within the FNO. We employ a linear range normalisation scheme, placing the field values between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser (Kingma and Ba, 2015) with a step decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a quantile loss in the case of CQR and an MSE loss in the other cases.
Figure 24: Waves: Temporal evolution of field associated with the wave equation modelled using the numerical spectral solver (top of the figure) and that of the U-Net (bottom of the figure). The spatial domain is given in Cartesian geometry.
Figure 25: Waves: Temporal evolution of field associated with the wave equation modelled using the numerical spectral solver (top of the figure) and that of the FNO (bottom of the figure). The spatial domain is given in Cartesian geometry.
@@ D.1 Physics
The Navier-Stokes scenario that we are interested in modelling is taken from the exact formulation in Li et al. (2021), where the viscosity of the incompressible fluid in 2D is expressed as:
@@
where $u$ is the velocity field and vorticity is the curl of the velocity field $w=\nabla\crossproduct u$. The domain is split across the spatial domain characterised by $x,y$ and the temporal domain $t$. The initial vorticity is given by the field $w_{0}$. The forcing function is given by $f$ and is a function of the spatial domain in $x,y$. We utilise two datasets from Li et al. (2021) that are built by solving the above equations with viscosities $\nu=1e-3$ and $\nu=1e-4$ under different initial vorticity distributions. For further information on the physics and the data generation, refer Li et al. (2021).
@@ D.2 Model and Training
We train an FNO to map the spatio-temporal evolution of the vorticity taking in the first 10 time instances (T_in) to the next 10 time instances (step). For the case of the Navier-Stokes equations, we deploy a feed-forward mapping from the initial 10 time steps to the next 10 time steps. The architecture for the FNO can be found in Table 8. We deploy a Min-Max normalisation strategy, allowing the field values to lie between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser (Kingma and Ba, 2015) with a step decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a relative LP loss. Considering the efficiency and simplicity, conformal prediction for the Navier-Stokes case was conducted using AER and STD as a nonconformity metric as given in the Section 2.2.
@@ D.2.1 Prediction
Figure 26: Navier–Stokes: Temporal evolution of the vorticity associated with the Navier-Stokes equations. In (a) we compare the FNO performance against that of the numerical solver within the training distribution ($\nu=1e-3$). In (b) we demonstrate the performance of the same FNO on out-of-distribution data upon which we perform Conformal Prediction ($\nu=1e-4$) modelled using the numerical solver (top of the figure) and that of the FNO (bottom of the figure). The spatial domain is given in Cartesian geometry.
@@ E.1 Physics
Within each simulation, multiple-density blobs with varying positions, width and amplitude are initialised in a low-density background. In the absence of a plasma current to hold the density blob in place, the pressure gradient term in the momentum equation generates a buoyancy effect, causing the blob to move outwards. The system under consideration is characterised by a highly correlated multi-variable setting as given above. Within each simulation, we evolve the blobs migrate radially outward until they reach the wall, where the Dirichlet boundary conditions engage to allow for convection and diffusion. Refer to (Gopakumar et al., 2024) for more detailed information about the setup.
@@
Z - Position of Blobs $U[-0.4,+0.4]$ Continuous
@@ E.2 Model and Training
We train a multi-variable FNO to map the spatio-temporal evolution of the field variable taking in the first 10 time instances (T_in) to the next 5 time instances (step). For the case of the MHD equations, we deploy an auto-regressive structure that performs a time rollout allowing us to map the initial time steps in a recursive manner up until the desired time instance (T_out). Each model autoregressively models the evolution of the field variable up until the $50^{th}$ time instance. The architecture for the multi-variable FNO can be found in Table 10. We deploy a two-fold normalisation strategy considering the nature of the dataset. The physical field information represented within the MHD cases are in different scales, with densities ranging from 0 to $201e$ 20 and temperatures ranging up to $61e$ 6. Since we are considering the gradual diffusion of an inhomogeneous density blob(s), the data distribution within the spatial domain is severely imbalanced. Taking these aspects of the training data into consideration, a physics normalisation is performed initially, where the field values are scaled down by dividing against the prominent field value. This is followed up by a linear range scaling, allowing the field values to lie between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser (Kingma and Ba, 2015) with a step decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a relative LP loss. Considering the efficiency and simplicity, conformal prediction for the MHD case was only conducted using the AER as a nonconformity metric as given in the Section 2.2.
@@ Appendix G Impact of Calibration Size
Across almost all the experiments in Section 3, we use a calibration dataset of the size 1000 exchangeable simulations. The only exception is for the multi-variable FNO for MHD, for which we use 100 simulations as there was limited data available. We chose the number 1000 as the baseline size of the calibration dataset since in (Angelopoulos and Bates, 2023) they demonstrate that choosing $n_{cal}=1000$ calibration points leads to a coverage that is typically between 0.88 and 0.92 for $\alpha=0.1$. Since the size and nature of the calibration set is a source of finite sample variability, it requires analysis across each problem to which we deploy CP.
Ideally Equation 1 holds for a calibration dataset of any size $n_{cal}$. The coverage guaranteed by CP conditionally on this calibration dataset is essentially a random quantity. Thus, depending on the choice of the calibration dataset, the coverage would fluctuate around $1-\alpha$. The distribution of the coverage as a function of the size of the calibration size is governed by a Beta distribution as given in Equation 4.
@@
Figure 30: Impact of the size of the calibration dataset on the coverage obtained for Conformal Prediction across various nonconformity scores for the Convection-Diffusion Equation. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage, however with larger $n_{cal}$ we obtain marginally better coverage. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage.
@@
From Figure 29 to Figure 32, we explore the impact of the size of the calibration dataset on the coverage obtained for the various nonconformity scores. We notice that across our experiments, we obtain guaranteed coverage, irrespective of the chosen nonconformity score or the size of the calibration dataset ($n_{cal}$). Though the size of the calibration dataset has no impact on the guarantee as can be witnessed in the above figures, the tightness of the error bars is governed by the size of the calibration dataset.
Appendix H Additional Results for Weather Forecasting Models
We here include some additional plots from the weather forecasting experiments. In Figure 33 we show slice plots along the limited area spatial domain. This further illustrates the error bars obtained at various coverage levels using the CP framework. Figure 34 shows empirical coverage for example variables in global weather forecasting.
@@ Appendix H Additional Results for Weather Forecasting Models
Figure 33: Slice plots across the $x$-axis of a temporal prediction of a single variable (u-component of wind). Figures (a) - (d) depicts the ground truth, prediction, upper and lower bars obtained through the CP framework for the Graph-FM (MSE) and Graph-FM (NLL) for 95 percent coverage ($\alpha=0.05$) and 15 percent coverage ($\alpha=0.85$).
@@
Figure 34: Empirical coverage for global weather forecasting models. We here plot the empirical coverage per variable, for temperature (t700) and geopotential (q700) at 700 hPa.

### v3

@@ C.2 Model and Training
We train U-Nets and FNOs to map the spatio-temporal evolution of the field variables. For the U-Nets, the network takes in the first 20 time instances (T_in) to map the next 30 time instances (step). The U-net performs a feed-forward mapping without any autoregressive roll-outs. For the FNO we deploy an auto-regressive structure that performs time rollouts, allowing us to map the initial time steps in a recursive manner up until the desired time instance (T_out). Each model autoregressively models the evolution of the field variable up until the $80^{th}$ time instance. The U-Net architecture can be found in Table 6 and the FNO in Table 7. For the case of STD, the architecture is modified with 2D dropout layers following each encoder and decoder of the U-Net and after each Fourier layer within the FNO. We employ a linear range normalisation scheme, placing the field values between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser (adam) with a step decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a quantile loss in the case of CQR and an MSE loss in the other cases.
Figure 24: Waves: Temporal evolution of the field associated with the wave equation modelled using the numerical spectral solver (top of the figure) and that of the U-Net (bottom of the figure). The spatial domain is given in Cartesian geometry.
Figure 25: Waves: Temporal evolution of the field associated with the wave equation modelled using the numerical spectral solver (top of the figure) and that of the FNO (bottom of the figure). The spatial domain is given in Cartesian geometry.
@@ D.1 Physics
The Navier-Stokes scenario that we are interested in modelling is taken from the exact formulation in li2021fourier, where the viscosity of the incompressible fluid in 2D is expressed as:
@@
where $u$ is the velocity field and vorticity is the curl of the velocity field $w=\nabla\crossproduct u$. The domain is split across the spatial domain characterised by $x,y$ and the temporal domain $t$. The initial vorticity is given by the field $w_{0}$. The forcing function is given by $f$ and is a function of the spatial domain in $x,y$. We utilise two datasets from li2021fourier that are built by solving the above equations with viscosities $\nu=1e-3$ and $\nu=1e-4$ under different initial vorticity distributions. For further information on the physics and the data generation, refer li2021fourier.
@@ D.2 Model and Training
We train an FNO to map the spatio-temporal evolution of the vorticity, taking in the first 10 time instances (T_in) to the next 10 time instances (step). For the case of the Navier-Stokes equations, we deploy a feed-forward mapping from the initial 10 time steps to the next 10 time steps. The architecture for the FNO can be found in Table 8. We deploy a Min-Max normalisation strategy, allowing the field values to lie between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser (adam) with a step decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a relative LP loss. Considering the efficiency and simplicity, conformal prediction for the Navier-Stokes case was conducted using AER and STD as a nonconformity metric as given in the Section 2.2.
@@ D.2.1 Prediction
Figure 26: Navier–Stokes: Temporal evolution of the vorticity associated with the Navier-Stokes equations. In (a), we compare the FNO performance against that of the numerical solver within the training distribution ($\nu=1e-3$). In (b), we demonstrate the performance of the same FNO on out-of-distribution data upon which we perform Conformal Prediction ($\nu=1e-4$) modelled using the numerical solver (top of the figure) and that of the FNO (bottom of the figure). The spatial domain is given in Cartesian geometry.
@@ E.1 Physics
Within each simulation, multiple-density blobs with varying positions, width and amplitude are initialised in a low-density background. In the absence of a plasma current to hold the density blob in place, the pressure gradient term in the momentum equation generates a buoyancy effect, causing the blob to move outwards. The system under consideration is characterised by a highly correlated multi-variable setting as given above. Within each simulation, we evolve the blobs to migrate radially outward until they reach the wall, where the Dirichlet boundary conditions engage to allow for convection and diffusion. Refer to (Gopakumar_2024) for more detailed information about the setup.
@@
Z - Position of Blobs U[0.4,+0.4]U[-0.4,+0.4] Continuous
@@ E.2 Model and Training
We train a multi-variable FNO to map the spatio-temporal evolution of the field variable, taking in the first 10 time instances (T_in) to the next 5 time instances (step). For the case of the MHD equations, we deploy an auto-regressive structure that performs a time rollout, allowing us to map the initial time steps in a recursive manner up until the desired time instance (T_out). Each model autoregressively models the evolution of the field variable up until the $50^{th}$ time instance. The architecture for the multi-variable FNO can be found in Table 10. We deploy a two-fold normalisation strategy considering the nature of the dataset. The physical field information represented within the MHD cases is in different scales, with densities ranging from 0 to $1e20$ and temperatures ranging up to $1e6$. Since we are considering the gradual diffusion of an inhomogeneous density blob(s), the data distribution within the spatial domain is severely imbalanced. Taking these aspects of the training data into consideration, a physics normalisation is performed initially, where the field values are scaled down by dividing by the prominent field value. This is followed up by a linear range scaling, allowing the field values to lie between -1 and 1. Each model is trained for up to 500 epochs using the Adam optimiser (adam) with a step decaying learning rate. The learning rate is initially set to 0.005 and scheduled to decrease by half after every 100 epochs. The model was trained using a relative LP loss. Considering the efficiency and simplicity, conformal prediction for the MHD case was only conducted using the AER as a nonconformity metric, as given in the Section 2.2.
@@ Appendix G Impact of Calibration Size
Across almost all the experiments in Section 3, we use a calibration dataset of the size 1000 exchangeable simulations. The only exception is for the multi-variable FNO for MHD, for which we use 100 simulations, as there was limited data available. We chose the number 1000 as the baseline size of the calibration dataset since in (gentle_introduction_CP) they demonstrate that choosing $n_{cal}=1000$ calibration points leads to a coverage that is typically between 0.88 and 0.92 for $\alpha=0.1$. Since the size and nature of the calibration set are a source of finite sample variability, it requires analysis across each problem to which we deploy CP.
Ideally, Equation 1 holds for a calibration dataset of any size $n_{cal}$. The coverage guaranteed by CP conditionally on this calibration dataset is essentially a random quantity. Thus, depending on the choice of the calibration dataset, the coverage would fluctuate around $1-\alpha$. The distribution of the coverage as a function of the size of the calibration size is governed by a Beta distribution as given in Equation 4.
@@
Figure 30: Impact of the size of the calibration dataset on the coverage obtained for Conformal Prediction across various nonconformity scores for the Convection-Diffusion Equation. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage; however, with larger $n_{cal}$ we obtain marginally better coverage. Irrespective of the chosen size of the calibration dataset, we obtain guaranteed coverage.
@@
From Figure 29 to Figure 32, we explore the impact of the size of the calibration dataset on the coverage obtained for the various nonconformity scores. We notice that across our experiments, we obtain guaranteed coverage, irrespective of the chosen nonconformity score or the size of the calibration dataset ($n_{cal}$). Though the size of the calibration dataset has no impact on the guarantee, as can be witnessed in the above figures, the tightness of the error bars is governed by the size of the calibration dataset.
Appendix H Neural Weather Prediction Models
H.1 Physics and Data
H.1.1 Limited Area Forecasting - MEPS Dataset
The MetCoOp Ensemble Prediction System (MEPS) provides operational weather forecasts for the Nordic region (meps). Our dataset comprises historical MEPS forecasts from April 2021 to March 2023, initialised at 00 and 12 UTC daily with 5 ensemble members per initialisation (sim\ sim10 forecasts per day).
Spatial Configuration:
The domain uses a Lambert conformal conic projection, downsampled from 2.5 km to 10 km resolution for computational efficiency, yielding a grid of $N_{x}\times N_{y}=238\times 268$ nodes.
Temporal Configuration:
Forecasts extend 66 hours with 1-hour steps. Models predict at 3-hour intervals using the last two states as input, resulting in an effective 57-hour evaluation horizon (19 time steps).
Variables:
The dataset includes $N_{\text{var}}=17$ variables: surface conditions (ground/sea level pressure, 2m temperature and humidity), radiation fluxes (net longwave/shortwave), lowest atmospheric level ( 12.5m: temperature, humidity, wind components), pressure levels (temperature at 500/850 hPa, wind at 850 hPa, geopotential at 500/1000 hPa), and integrated water vapor column. All variables except solar radiation are instantaneous.
Data Splits:
Training uses months 1-15 (April 2021-June 2022), validation uses months 16-18, calibration uses September 2021 forecasts (from 2021-09-04 onwards), and testing uses September 2022 forecasts. Using the same month for calibration and testing minimises seasonal distributional shifts.
H.1.2 Global Forecasting - ERA5 Dataset
Global models are trained on ERA5 reanalysis data (era5) at 1.5° resolution ($N_{x}\times N_{y}=240\times 121$ nodes). Forecasts extend 10 days with 6-hour steps ($T=40$ time steps). The full dataset contains 83 variables (5 surface variables and 6 atmospheric variables at 13 pressure levels). For computational efficiency, CP experiments use $N_{\text{var}}=11$ variables: all 5 surface variables and the 6 atmospheric variables at 700 hPa. Training uses multiple years of historical data, calibration uses the full year 2018, and testing uses 2019.
H.2 Model Architecture and Training
H.2.1 Graph Neural Network Architecture
The Graph-FM architecture (neural_lam) employs a hierarchical GNN with an encode-process-decode structure. Grid nodes represent the spatial discretisation, while mesh nodes form a coarser multi-resolution hierarchy ($L=4$ levels for a limited area, $L=8$ for global). Encoding maps grid states to the mesh, processing performs message passing across mesh levels (4 or 8 layers), and decoding projects back to the grid. All representations use 64-dimensional embeddings (limited area) or 256-dimensional embeddings (global).
H.2.2 Training Procedure
Models undergo two-stage training: (1) single-step prediction ($X^{t+1}$ from $X^{t-1:t}$), then (2) rollout fine-tuning with 4-step autoregressive sequences. Two model variants are trained:
Graph-FM (MSE): Uses weighted MSE loss for deterministic predictions:
$\mathcal{L}_{\text{MSE}}=\frac{1}{N_{\text{rollout}}}\sum_{t}\frac{1}{|G|}\sum_{v\in G}\sum_{i=1}^{N_{\text{var}}}\lambda_{i}\omega_{i}\left(\hat{X}^{t}_{v,i}-X^{t}_{v,i}\right)^{2}$ (19)
where $\lambda_{i}$ is the inverse variance of time differences, $\omega_{i}$ weights the vertical level, and $G$ is the set of non-boundary grid nodes.
Graph-FM (NLL): Uses NLL loss with diagonal Gaussians, outputting mean $\mu$ and standard deviation $\sigma$:
$\mathcal{L}_{\text{NLL}}=\frac{1}{N_{\text{rollout}}}\sum_{t}\frac{1}{|G|}\sum_{v\in G}\sum_{i=1}^{N_{\text{var}}}\left[\log\sigma^{t}_{v,i}+\frac{(X^{t}_{v,i}-\mu^{t}_{v,i})^{2}}{2(\sigma^{t}_{v,i})^{2}}\right]$ (20)
Training uses AdamW (learning rate 0.001), batch size 8 (limited area), 500 epochs for single-step and 200 for rollout fine-tuning, requiring 3-4 days on an NVIDIA A100 GPU (80 GB). All variables are normalised to zero mean and unit variance using training statistics.
H.3 Inference and Autoregressive Forecasting
Given initial states $X^{-1:0}$ and forcing inputs $F^{1:T}$, models autoregressively predict $X^{t+1}=\hat{f}(X^{t-1:t},F^{t+1})$. Limited area models use lateral boundary forcing: predictions within a 10-grid-cell boundary are replaced with ground truth at each step. Forcing inputs include solar radiation at the top-of-atmosphere, diurnal and annual cycle encodings (sine/cosine transforms), and open water fraction. A 57-hour limited area forecast requires only 1.5 seconds on a single A100 GPU.
H.4 Conformal Prediction Application
H.4.1 Nonconformity Scores
For Graph-FM (MSE) (Absolute Error Residual):
$s(X,Y)=|Y-\hat{f}(X)|,\quad\mathbb{C}^{\alpha}(X)=[\hat{f}(X)-\hat{q},\hat{f}(X)+\hat{q}$] (21)
For Graph-FM (NLL) (Standard Deviation):
$s(X,Y)=\frac{|Y-\mu(X)|}{\sigma(X)},\quad\mathbb{C}^{\alpha}(X)=[\mu(X)-\hat{q}\cdot\sigma(X),\mu(X)+\hat{q}\cdot\sigma(X)]$ (22)
This calibrates the model’s potentially miscalibrated uncertainty estimates.
H.4.2 Calibration Procedure
CP is performed independently for each spatio-temporal cell. Limited area models calibrate $19\times 238\times 268\times 17=20{,}602{,}232$ cells using  270 September 2021 forecasts; global models use  730 forecasts from 2018. For each cell, the $(1-\alpha)$-quantile is computed:
$\hat{q}=F^{-1}_{\hat{s}}\left(\frac{\lceil(n+1)(1-\alpha)\rceil}{n}\right)$ (23)
Calibration is computationally efficient: 229-310 seconds for limited area models and 366-401 seconds for global models, including score computation, quantile calculation, and coverage validation.
H.5 Exchangeability Considerations
H.5.1 Weather Forecasting as an Initial Boundary Value Problem
Applying CP to time series typically violates exchangeability due to temporal dependencies. However, treating each forecast as an independent Initial Boundary Value Problem (IBVP) preserves exchangeability: each forecast is fully determined by its initial conditions $X^{-1:0}$ and forcing $F^{1:T}$, making forecasts initialised at sufficiently separated times independent realisations from the atmospheric state space. This holds even when forecasts cover overlapping time periods, since each prediction conditions only on its own initial state.
CP is applied to the entire spatio-temporal output tensor $Y\in\mathbb{R}^{T_{\text{out}}\times N_{x}\times N_{y}\times N_{\text{var}}}$ simultaneously, preserving the physical forecast structure.
H.5.2 Calibration-to-Test Exchangeability
The coverage guarantee requires test forecasts to be exchangeable with calibration forecasts. For limited area models, this assumes September 2022 patterns are exchangeable with September 2021 patterns; for global models, 2019 states are exchangeable with 2018 states. These assumptions are reasonable given year-to-year climate consistency, validated by the excellent empirical coverage observed. Potential violations include anomalous events not represented in calibration data, systematic climate shifts, or changes in model distribution, warranting periodic recalibration with recent data.
To investigate the setting where there is a distributional shift between the calibration and test datasets, we consider a setting where we shift the month of the year in the weather data. For this experiment, we calibrate the limited area weather model using predictions for September 2021, and evaluate on January 2023. The weather in the Nordic region being modelled exhibits strong seasonal dependencies, with the winter month of January being much colder than the September month used for calibration. Note that while we only use September for calibration, the model has been trained with data from all seasons of the year. Empirical coverage is shown in Figure 33, and in Figure 34 we plot the width of predictive intervals for 2 m temperature predictions. We note that, while the coverage shows larger deviations than in the setting where both calibration and evaluation were performed for the September data, the method overall still provides useful uncertainty estimates. This may be attributed to the reason that the physics regime characterising the weather evolution does not change across the months, but only the initial conditions to which the model is exposed.
(a) Graph-FM (MSE)
(b) Graph-FM (NLL)
Figure 33: Empirical Coverage for limited area weather forecasting models, when calibrated on September and evaluating coverage on January.
(a) Graph-FM (MSE)
(b) Graph-FM (NLL)
Figure 34: Width of predictive intervals in an example January forecast from the distributional shift setting. Here $\alpha=0.05$ and we show results for temperature 2 m above ground (2t).
H.6 Additional Results
Figure 35 shows spatial slices illustrating error bars at different coverage levels. The Graph-FM (NLL) model produces tighter, spatially variable error bars due to input-dependent uncertainties, while Graph-FM (MSE) produces constant-width bars at each lead time. Both achieve target coverage, validating the CP framework.
Figure 36 demonstrates empirical coverage for global forecasting (temperature and geopotential at 700 hPa). Both model types closely follow the ideal diagonal, confirming guaranteed marginal coverage across 1,161,6001{,}161{,}600 spatial-temporal points per variable.
H.6.1 Deterministic vs. Probabilistic Models
Deterministic Models (AER): Error bars have constant width at each lead time, determined purely by calibration performance. Computationally efficient but potentially overconservative in predictable regions.
Probabilistic Models (STD): Error bars adapt to local forecast uncertainty, providing tighter bounds in high-confidence regions. Requires NLL training but benefits from CP calibration of potentially miscalibrated model uncertainties.
Both approaches achieve guaranteed coverage, with probabilistic models generally providing more informative bounds at the cost of additional complexity.
@@ H.6 Additional Results
Figure 35: Slice plots across the $x$-axis showing u-wind component predictions. Panels (a)-(d) show ground truth, prediction, and CP-derived upper/lower bounds for Graph-FM (MSE) and Graph-FM (NLL) at 95% ($\alpha=0.05$) and 15% ($\alpha=0.85$) coverage.
@@
Figure 36: Empirical coverage for global weather forecasting models showing temperature (t700) and geopotential (q700) at 700 hPa. Both models achieve near-ideal coverage across all $\alpha$ levels.